OpenStudy (danieke):

HELP PLEASE A car is traveling at a speed of 81 miles per hour. What is the car's speed in feet per second? How many feet will the car travel in 15 seconds? In your computations, use the fact that 1 mile is equal to 5280 feet. Do not round your answers. speed ft/s: distance traveled in 15 seconds ft:

OpenStudy (anonymous):

How Many minutes are in an hour? How many Seconds are in an hour?

OpenStudy (danieke):

60 and 3600

OpenStudy (anonymous):

3600/18 = ?

OpenStudy (anonymous):

Let me see your options as well. Is this multiple choice?

OpenStudy (danieke):

no its free response

OpenStudy (anonymous):

okay like she said what is 3600 divided by 18=?

OpenStudy (danieke):

200

OpenStudy (anonymous):

look im very sorry but i have a dear friends that would love to help i have school in the morning @zepdrix trust me he is smarter then me and btw he dosent like ppl who just want the answer soo work for it im sorry @zepdrix will help

zepdrix (zepdrix):

\[\large\rm \frac{81miles}{hour}\]So umm :) what were you doing? Going from hours to minutes? Then minutes to seconds I guess? :D

OpenStudy (danieke):

81/over?

zepdrix (zepdrix):

\[\large\rm \frac{81miles}{hour}\cdot\frac{1hour}{60minutes}\]We would like hours to cancel out, so we make sure it's in the top for this new conversion. This is a unit, 1hour is equivalent to 60 minutes. So we're allowed to do this.

zepdrix (zepdrix):

What's up? :) Is that confusing why the 81 is in the numerator?

OpenStudy (danieke):

yes it is

zepdrix (zepdrix):

No no no, let's leave it in the numerator :D It'll get confusing otherwise.

zepdrix (zepdrix):

The 81 is the `number of miles` for every hour traveled. The number is connected to the numerator.

zepdrix (zepdrix):

Try to make your peace with it XD lol

OpenStudy (danieke):

okay i got that part

zepdrix (zepdrix):

This is where we want to apply "cross cancellation". We have hours in our numerator AND denominator, so the units will cancel out.\[\large\rm \frac{81miles}{\cancel{hour}}\cdot\frac{\cancel{hour}}{60minutes}\]

OpenStudy (danieke):

27/20?

zepdrix (zepdrix):

Oh simplifying? Sure sure sure :)\[\large\rm \frac{27miles}{20minutes}\]We want to go from minutes to seconds. Again, since our minutes are in the `denominator`, we would like the minutes to show up in the `numerator` of our unit fraction,

zepdrix (zepdrix):

\[\large\rm \frac{27miles}{20minutes}\cdot \frac{1minute}{60seconds}\]

zepdrix (zepdrix):

\[\large\rm \frac{27miles}{20\cancel{minutes}}\cdot \frac{1\cancel{minute}}{60seconds}\]

zepdrix (zepdrix):

\[\large\rm \frac{27miles}{20\cdot60seconds}\]

OpenStudy (danieke):

cross multiply?

zepdrix (zepdrix):

With fractions, we always multiply `straight across`. You can cross-multiply when there is an equality sign between them. But there is no equals sign here, so try to get cross-multiply out of your head ^^

OpenStudy (danieke):

9/400?

zepdrix (zepdrix):

\[\large\rm \frac{9miles}{400seconds}\]Ok great. Notice our denominator is perfect now. We're trying to get to `feet` / `second` We have the seconds for our units in the bottom.

zepdrix (zepdrix):

Now we'll convert `miles` to `feet` using the conversion factor that they gave us: 1mile = 5280feet

zepdrix (zepdrix):

Notice that miles is in the numerator, so for our conversion fraction, we want to put miles in the denominator so we can get the cancellation we're looking for.\[\large\rm \frac{9miles}{400seconds}\cdot\frac{5280feet}{1mile}\]

OpenStudy (danieke):

so cancel this?

zepdrix (zepdrix):

Yes, cancel "miles", and then let's see if we can simplify anywhere.

zepdrix (zepdrix):

\[\large\rm \frac{9\cancel{miles}}{400seconds}\cdot\frac{5280feet}{1\cancel{mile}}\]

OpenStudy (danieke):

594/5

OpenStudy (danieke):

?

zepdrix (zepdrix):

Ok great. So we have \(\large\rm \dfrac{594feet}{5seconds}\) 594 feet in 5 seconds. Another way we can write this (which will look better for your answer): \(\large\rm \dfrac{594}{5}\dfrac{feet}{sec}\) (594/5) feet per second.

zepdrix (zepdrix):

So there is our speed :) 594/5 ft/s

zepdrix (zepdrix):

This is the distance traveled EVERY SECOND. So if we want the distance traveled in 15 seconds, we want 15 of these fractions. So we simply multiply our speed by (15 seconds)

OpenStudy (danieke):

294/5 *15

zepdrix (zepdrix):

\[\large\rm \dfrac{594feet}{5seconds}\cdot 15seconds\]Think of these new seconds as being in the "numerator", so again we have cancellation.\[\large\rm \dfrac{594feet}{5\cancel{seconds}}\cdot 15\cancel{seconds}\]Good, yes. And make sure you put the 15 in the numerator, or in other words, multiply the 15 and 294, he doesn't go in the bottom.

zepdrix (zepdrix):

woops it was a 594 right? :) Not 294 hehe

OpenStudy (danieke):

yes it was ;)

zepdrix (zepdrix):

And then ya, simplify if you're able.

OpenStudy (danieke):

198?

zepdrix (zepdrix):

Hmm that seems way too small :o

zepdrix (zepdrix):

15 * 594 / 5 = ?

OpenStudy (danieke):

1782?

OpenStudy (danieke):

am i right ?

zepdrix (zepdrix):

1782 feet. Yes good job \c:/

OpenStudy (danieke):

thank you so much @zepdrix